The marketing manager of a department store has determined that revenue, in dollars, is related to the number of units of television advertising, x, and the number of units of newspaper advertising, y, by the function R(x,y)=850(87x−2y^2+3xy−3x^2). Each unit of television advertising costs $1000, and each unit of newspaper advertising costs $500. If the amount spent on advertising is $11500, find the maximum revenue.
Answer #1

The marketing manager of a department store has determined that revenue, in dollars, is related to the number of units of television advertising, x, and the number of units of newspaper advertising, y, by the function R(x,y)=850(87x−2y^2+3xy−3x^2). Each unit of television advertising costs $1000, and each unit of newspaper advertising costs $500. If the amount spent on advertising is $11500, find the maximum revenue.

solutions Given Roy) 850 (878-9y2 + 3xy –3x2) No. of units of television Advertising go no. of units of newspapers AdvertisinNow 6 84 6x = 8y =) <= 8(8715) =) x i) 6 116 -) ca 5 0=(Rxa)(Ryy) -(Rxy)! (6*850) (-4*850) - (850*3)² D = 6 = 850?(24-9) - 85